OpenStudy (anonymous):

factor the expression and solve it, plz show work. 18x^2+48x=-32

OpenStudy (anonymous):

well first go 18 times 18 then combine like terms

OpenStudy (anonymous):

1) First add 32 to both sides to make the equation 18x^2+48x+32=0 2) Factor out a 2 to get 2(9x^2+24x+16)=0 3) Factor 9x^2+24x+16 and you get (3x+4)(3x+4) ... FOIL it to make sure 4) Then you equation is 2(3x+4)(3x+4)=0 5) Solve for x and you get 1 solution x= -4/3 since it is a perfect square Hope that helps :D